# Vanilla Transformers are Transfer Capability Teachers

### v1

@@ Abstract
Vanilla Transformers are Transfer Capability Teachers
Xin Lu, Yanyan Zhao††thanks: Email corresponding., Bing Qin Research Center for Social Computing and Information Retrieval Harbin Institute of Technology, China {xlu, yyzhao, qinb}@ir.hit.edu.cn
@@ 2.3 Transfer Capability Distillation
Model | Pre-Train | Pre-Train | CoLA | MRPC | SST-2 | STS-B | RTE | MNLI | QNLI | QQP | Avg. | |||||
Epoch | Pref. | (8.5k) | (3.7k) | (67k) | (7.0k) | (2.5k) | (393k) | (108k) | (364k) | Score | ||||||
Pre-Training Performance Alignment (H=128) | ||||||||||||||||
Vanilla BERT (Teacher) | 20.0 | -2.65387 | 33.88 | 83.03 | 88.09 | 86.65 | 83.44 | 83.39 | 63.41 | 76.77 | 77.36 | 85.15 | 88.71 | 84.82 | 77.89 | |
MoE BERT | 6.0 | -2.32278 | 37.56 | 82.11 | 87.15 | 86.22 | 83.68 | 83.31 | 62.86 | 74.94 | 76.13 | 84.97 | 87.87 | 83.90 | 77.56 | |
MoE BERT w/ TCD | 5.0 | -2.33650 | 44.10 | 84.52 | 89.09 | 87.41 | 84.08 | 83.81 | 65.70 | 77.27 | 78.29 | 86.12 | 88.72 | 85.01 | 79.51 | |
Pre-Training Performance Alignment (H=768) | ||||||||||||||||
Vanilla BERT (Teacher) | 10.0 | -1.54597 | 62.03 | 86.93 | 90.65 | 92.73 | 87.64 | 87.31 | 61.97 | 83.73 | 83.88 | 90.72 | 90.71 | 87.44 | 83.81 | |
MoE BERT | 12.0 | -1.29679 | 64.24 | 86.11 | 90.29 | 93.12 | 87.25 | 86.89 | 61.73 | 83.61 | 83.61 | 90.26 | 90.29 | 87.26 | 83.72 | |
MoE BERT w/ TCD | 10.0 | -1.30669 | 65.36 | 88.03 | 91.53 | 93.46 | 88.10 | 87.79 | 64.14 | 84.65 | 84.68 | 91.63 | 90.85 | 87.70 | 84.83 |
@@ 2.4 Training Process
Model | Teacher | Pre-Train | Pre-Train | CoLA | MRPC | SST-2 | STS-B | RTE | MNLI | QNLI | QQP | Avg. | |||||
Epoch | Epoch | Pref. | (8.5k) | (3.7k) | (67k) | (7.0k) | (2.5k) | (393k) | (108k) | (364k) | Score | ||||||
Pre-Training Epoch Alignment (H=128) | |||||||||||||||||
MoE BERT | - | 25.0 | -2.08134 | 42.72 | 82.75 | 87.57 | 87.64 | 84.10 | 84.02 | 62.98 | 76.46 | 77.47 | 86.07 | 88.17 | 84.26 | 78.68 | |
MoE BERT w/ TCD | 20.0 | 5.0 | -2.33650 | 44.10 | 84.52 | 89.09 | 87.41 | 84.08 | 83.81 | 65.70 | 77.27 | 78.29 | 86.12 | 88.72 | 85.01 | 79.51 | |
Pre-Training Epoch Alignment (H=768) | |||||||||||||||||
MoE BERT | - | 20.0 | -1.20991 | 64.81 | 86.51 | 90.52 | 93.77 | 87.46 | 87.07 | 62.09 | 84.17 | 84.13 | 90.43 | 90.57 | 87.45 | 84.08 | |
MoE BERT w/ TCD | 10.0 | 10.0 | -1.30669 | 65.36 | 88.03 | 91.53 | 93.46 | 88.10 | 87.79 | 64.14 | 84.65 | 84.68 | 91.63 | 90.85 | 87.70 | 84.83 |
@@ 3.2.1 Pre-training Performance Alignment
Model | MRPC | STS-B | MNLI | QNLI | ||||
(H=128, Pre-Training Performance –2.56) | ||||||||
MoE BERT | 79.75 | 85.75 | 81.83 | 81.56 | 71.33 | 72.82 | 83.53 | |
MoE BERT + T | 80.33 | 86.05 | 82.53 | 82.21 | 74.77 | 75.72 | 84.61 | |
MoE BERT + T, I | 83.70 | 88.42 | 83.07 | 82.80 | 76.37 | 77.14 | 85.49 | |
MoE BERT + T, I, A | 82.87 | 88.02 | 82.96 | 82.74 | 75.45 | 76.25 | 85.40 | |
(H=768, Pre-Training Performance –1.42) | ||||||||
MoE BERT + T, I | 86.60 | 90.53 | 86.78 | 86.57 | 83.16 | 83.46 | 90.33 | |
MoE BERT + T, I, A | 87.58 | 91.28 | 87.38 | 87.10 | 83.48 | 83.70 | 90.80 |
@@ 6 Related Work
Model | AX | DM | GH | SE | UI | |
(H=128, Pre-Training Performance –2.76) | ||||||
Vanilla BERT | -3.545 | -2.955 | -3.462 | -3.530 | -4.120 | |
MoE BERT | -3.613 | -3.026 | -3.564 | -3.585 | -4.164 | |
MoE BERT w/ TCD | -3.563 | -2.959 | -3.499 | -3.536 | -4.118 | |
(H=768, Pre-Training Performance –1.57) | ||||||
Vanilla BERT | -2.338 | -2.179 | -2.420 | -2.443 | -3.052 | |
MoE BERT | -2.393 | -2.296 | -2.481 | -2.481 | -3.121 | |
MoE BERT w/ TCD | -2.334 | -2.219 | -2.426 | -2.444 | -3.051 |

### v2

# Improving the Downstream Performance of Mixture-of-Experts Transformers via Weak Vanilla Transformers

@@ Abstract
Improving the Downstream Performance of Mixture-of-Experts Transformers via Weak Vanilla Transformers
Xin Lu, Yanyan Zhao††thanks: Email corresponding., Bing Qin, Ting Liu Research Center for Social Computing and Interactive Robotics Harbin Institute of Technology, China {xlu, yyzhao, qinb, tliu}@ir.hit.edu.cn
@@ 2.3 Transfer Capability Distillation
| Model | Pre-Train | Pre-Train | CoLA | MRPC | SST-2 | STS-B | RTE | MNLI | QNLI | QQP | Avg. | ||||||
| Epoch | Pref. | (8.5k) | (3.7k) | (67k) | (7.0k) | (2.5k) | (393k) | (108k) | (364k) | Score | |||||||
| Pre-Training Performance Alignment (H=128) | |||||||||||||||||
| Vanilla BERT (Teacher) | 20.0 | -2.65387 | 33.88 | 83.03 | 88.09 | 86.65 | 83.44 | 83.39 | 63.41 | 76.77 | 77.36 | 85.15 | 88.71 | 84.82 | 77.89 | ||
| MoE BERT | 6.0 | -2.32278 | 37.56 | 82.11 | 87.15 | 86.22 | 83.68 | 83.31 | 62.86 | 74.94 | 76.13 | 84.97 | 87.87 | 83.90 | 77.56 | ||
| MoE BERT w/ TCD | 5.0 | -2.33650 | 44.10 | 84.52 | 89.09 | 87.41 | 84.08 | 83.81 | 65.70 | 77.27 | 78.29 | 86.12 | 88.72 | 85.01 | 79.51 | ||
| Pre-Training Performance Alignment (H=768) | |||||||||||||||||
| Vanilla BERT (Teacher) | 10.0 | -1.54597 | 62.03 | 86.93 | 90.65 | 92.73 | 87.64 | 87.31 | 61.97 | 83.73 | 83.88 | 90.72 | 90.71 | 87.44 | 83.81 | ||
| MoE BERT | 12.0 | -1.29679 | 64.24 | 86.11 | 90.29 | 93.12 | 87.25 | 86.89 | 61.73 | 83.61 | 83.61 | 90.26 | 90.29 | 87.26 | 83.72 | ||
| MoE BERT w/ TCD | 10.0 | -1.30669 | 65.36 | 88.03 | 91.53 | 93.46 | 88.10 | 87.79 | 64.14 | 84.65 | 84.68 | 91.63 | 90.85 | 87.70 | 84.83 | ||
@@ 2.4 Training Process
| Model | Teacher | Pre-Train | Pre-Train | CoLA | MRPC | SST-2 | STS-B | RTE | MNLI | QNLI | QQP | Avg. | ||||||
| Epoch | Epoch | Pref. | (8.5k) | (3.7k) | (67k) | (7.0k) | (2.5k) | (393k) | (108k) | (364k) | Score | |||||||
| Pre-Training Epoch Alignment (H=128) | ||||||||||||||||||
| MoE BERT | - | 25.0 | -2.08134 | 42.72 | 82.75 | 87.57 | 87.64 | 84.10 | 84.02 | 62.98 | 76.46 | 77.47 | 86.07 | 88.17 | 84.26 | 78.68 | ||
| MoE BERT w/ TCD | 20.0 | 5.0 | -2.33650 | 44.10 | 84.52 | 89.09 | 87.41 | 84.08 | 83.81 | 65.70 | 77.27 | 78.29 | 86.12 | 88.72 | 85.01 | 79.51 | ||
| Pre-Training Epoch Alignment (H=768) | ||||||||||||||||||
| MoE BERT | - | 20.0 | -1.20991 | 64.81 | 86.51 | 90.52 | 93.77 | 87.46 | 87.07 | 62.09 | 84.17 | 84.13 | 90.43 | 90.57 | 87.45 | 84.08 | ||
| MoE BERT w/ TCD | 10.0 | 10.0 | -1.30669 | 65.36 | 88.03 | 91.53 | 93.46 | 88.10 | 87.79 | 64.14 | 84.65 | 84.68 | 91.63 | 90.85 | 87.70 | 84.83 | ||
@@ 3.2.1 Pre-training Performance Alignment
| Model | MRPC | STS-B | MNLI | QNLI | |||||
| (H=128, Pre-Training Performance –2.56) | |||||||||
| MoE BERT | 79.75 | 85.75 | 81.83 | 81.56 | 71.33 | 72.82 | 83.53 | ||
| MoE BERT + T | 80.33 | 86.05 | 82.53 | 82.21 | 74.77 | 75.72 | 84.61 | ||
| MoE BERT + T, I | 83.70 | 88.42 | 83.07 | 82.80 | 76.37 | 77.14 | 85.49 | ||
| MoE BERT + T, I, A | 82.87 | 88.02 | 82.96 | 82.74 | 75.45 | 76.25 | 85.40 | ||
| (H=768, Pre-Training Performance –1.42) | |||||||||
| MoE BERT + T, I | 86.60 | 90.53 | 86.78 | 86.57 | 83.16 | 83.46 | 90.33 | ||
| MoE BERT + T, I, A | 87.58 | 91.28 | 87.38 | 87.10 | 83.48 | 83.70 | 90.80 | ||
@@ 6 Related Work
| Model | AX | DM | GH | SE | UI | ||
| (H=128, Pre-Training Performance –2.76) | |||||||
| Vanilla BERT | -3.545 | -2.955 | -3.462 | -3.530 | -4.120 | ||
| MoE BERT | -3.613 | -3.026 | -3.564 | -3.585 | -4.164 | ||
| MoE BERT w/ TCD | -3.563 | -2.959 | -3.499 | -3.536 | -4.118 | ||
| (H=768, Pre-Training Performance –1.57) | |||||||
| Vanilla BERT | -2.338 | -2.179 | -2.420 | -2.443 | -3.052 | ||
| MoE BERT | -2.393 | -2.296 | -2.481 | -2.481 | -3.121 | ||
| MoE BERT w/ TCD | -2.334 | -2.219 | -2.426 | -2.444 | -3.051 | ||